# Compatibility of canonical $\ell$-adic local systems on Shimura varieties

### v1

@@ 3. Ramification
In [16, Proposition 2.3], we showed that Zariski-dense $G^{\mathrm{ad}}(\mathbb{C})$-valued representations $\rho$ of $\pi_{1}^{\mathrm{top}}(X(\mathbb{C}))$ in the “superrigid” (loc. cit. Definition 2.2) regime have arithmetic descents with independent of $\ell$ ramification (the same would work for Zariski-dense cohomologically rigid representations with quasi-unipotent local monodromy and finite-order abelianization). Without Zariski-density, and more precisely when the centralizer of $\rho$ is non-trivial, the argument breaks down, and indeed there is no reason that arbitrarily chosen arithmetic descents $\rho_{\lambda}\colon\pi_{1}(X_{F})\to G(\overline{\mathbb{Q}}_{\lambda})$ should have this property, since one can for varying $\lambda$ twist $\rho_{\lambda}$ by homomorphisms $\mathrm{Gal}(\overline{\mathbb{Q}}/F)\to Z_{G}(\rho_{\lambda})(\overline{\mathbb{Q}}_{\lambda})$ with different ramification sets (the basic case to keep in mind, even with $\rho^{\mathrm{ad}}$ Zariski-dense, is when this centralizer is the center of a non-adjoint group $G$). In this section (Lemma 3.2), we show that having one number field specialization with independent of $\ell$ ramification is enough to show that a collection of descents also has this property.
Let $G$ be a connected reductive group; since we will work with sufficiently large coefficients, we assume $G$ is split over $\mathbb{Z}$. In this section we work with a smooth quasi-projective variety $X$ over a number field $F$ whose topological fundamental group $\Gamma=\pi_{1}^{\mathrm{top}}(X(\mathbb{C}),\bar{x})$ (we will specify a base-point later) is $G^{\mathrm{ad}}$-superrigid in the sense of [16, Definition 2.2]: that is, for any algebraically closed field $\Omega$ of characteristic zero and any two Zariski-dense homomorphisms $\rho_{1},\rho_{2}\colon\Gamma\to G^{\mathrm{ad}}(\Omega)$, there is a $\tau\in\operatorname{Aut}(G^{\mathrm{ad}}_{\Omega})$ such that $\tau(\rho_{1})=\rho_{2}$. [^4] Alternatively, throughout this section we could take cohomologically rigid representations with bounded order of quasi-unipotency, without assuming superrigidity properties of $\Gamma$; somewhat longer versions of the arguments then apply using the technique of [8], [15]. We fix once and for all a smooth projective compactification $X\subset\overline{X}$ with $\overline{X}\setminus X$ a strict normal crossings divisor. We further let $N$ be a large enough integer that the setup $(X,\overline{X},\overline{X}\setminus X)$ spreads out to a relative strict normal crossings divisor over $\mathcal{O}_{F}[1/N]$, and we choose some $x\in X(\mathcal{O}_{F^{\prime}}[1/N])$ with $F^{\prime}/F$ a finite extension.
@@
Indeed, this follows from a standard argument (variant of [21, Theorem 4], [8], [15]): the outer action of $\mathrm{Gal}(\overline{\mathbb{Q}}/F)$ on $\pi_{1}(X_{\overline{\mathbb{Q}}},\bar{x})$ induces a discrete action on the finite set $\mathcal{R}_{Z}$, so there is an open subgroup $H\subset\mathrm{Gal}(\overline{\mathbb{Q}}/F)$ such that $H$ fixes each $\rho_{\lambda,\overline{\mathbb{Q}}}$ (up to isomorphism). The adjoint descent $\rho_{\lambda}^{\mathrm{ad}}$ exists uniquely (using Zariski-density: see [16, Lemma 2.1]) over $F(\lambda)^{\mathrm{ad}}:=\overline{\mathbb{Q}}^{H}$. As we vary $\rho$, the $\rho_{\lambda}^{\mathrm{ad}}$ all factor through $G^{\mathrm{ad}}(\mathcal{O})$ for $\mathcal{O}$ the ring of integers in a finite extension of $\mathbb{Q}_{\ell}$, and $G_{Z}(\mathcal{O})\to G^{\mathrm{ad}}(\mathcal{O})$ is on an open subgroup of the source an isomorphism onto its image (an open subgroup of the target). Therefore after restricting to a finite extension $F(\lambda)/F(\lambda)^{\mathrm{ad}}$, each $\rho_{\lambda}^{\mathrm{ad}}$ lifts to
@@
Each $\rho_{\lambda}$ factors through $\pi_{1}(X_{\mathcal{O}_{F(\lambda)}[1/N_{\lambda}]},\bar{x})$ for an integer $N_{\lambda}$ (which we may assume divisible by $\ell$) that depends on $\lambda$ but not on $\rho\in\mathcal{R}_{Z}$. As in [16, Proposition 2.3], set $N^{\prime}=\mathrm{gcd}(N_{\lambda})$, and suppose that $p$ does not divide $N^{\prime}$. We claim that for all $\lambda^{\prime}\nmid p$, $\rho_{\lambda^{\prime},\overline{\mathbb{Q}}}$ factors via the tame specialization map through $\pi_{1}^{t}(X_{\overline{\kappa(v)}},\bar{x}_{v})$ for every $v|p$. Suppose that $p$ divides $N_{\lambda^{\prime}}$ (else the assertion is clear). For some $\lambda\neq\lambda^{\prime}$, $p$ does not divide $N_{\lambda}$. Let $\bar{v}$ be any place of $\overline{\mathbb{Q}}$ above $p$, and let $v=\bar{v}|_{F(\lambda)}$, so for each $\rho\in\mathcal{R}_{Z}$ we have the (tame: see [16, Footnote 3]) restriction to the special fiber
@@
By [7, Theorem 1.4.1], $\rho_{\lambda,v}$ has a (strong) $\lambda^{\prime}$-companion
@@
which clearly lies in $\mathcal{R}_{Z}$. [^6] Had we not assumed superrigidity and instead assumed the elements of $\mathcal{R}_{Z}$ to be cohomologically rigid, this conclusion would still be valid but would require the arguments of [8] and [15]. We may view the elements of $\mathcal{R}_{Z}$ as ¯$\overline{\mathbb{Q}}_{\lambda^{\prime}}$-valued using that they all have representatives factoring through $G(\mathcal{O}_{L})\subset G(\overline{\mathbb{Q}}_{\lambda^{\prime}})$, and we claim that the map $\mathcal{R}_{Z}\to\mathcal{R}_{Z}$, $\rho\mapsto\rho_{\lambda\leadsto\lambda^{\prime},\bar{v}}^{\mathrm{top}}$, is a bijection. Indeed, let $\rho$ and $\sigma$ be elements of $\mathcal{R}_{Z}$ such that $\rho_{\lambda\leadsto\lambda^{\prime},\bar{v}}^{\mathrm{top}}=\sigma_{\lambda\leadsto\lambda^{\prime},\bar{v}}^{\mathrm{top}}$ (equality up to $G$-conjugacy). Then $\rho_{\lambda\leadsto\lambda^{\prime},\bar{v}}=\sigma_{\lambda\leadsto\lambda^{\prime}\bar{v}}$. Let $F_{v}\in\pi_{1}(X_{\kappa(v)},\bar{x}_{v})$ be the $v$-Frobenius induced by the section ($\kappa(v)$-rational point) $x_{v}$. The equality $\rho_{\lambda\leadsto\lambda^{\prime},\bar{v}}(F_{v}\gamma F_{v}^{-1})=\sigma_{\lambda\leadsto\lambda^{\prime},\bar{v}}(F_{v}\gamma F_{v}^{-1})$, for all $\gamma\in\pi_{1}(X_{\overline{\kappa(v)}},\bar{x}_{v})$ implies that $\sigma_{\lambda\leadsto\lambda^{\prime},v}(F_{v})^{-1}\rho_{\lambda\leadsto\lambda^{\prime},v}(F_{v})$ centralizes the image of $\rho_{\lambda\leadsto\lambda^{\prime},\bar{v}}=\sigma_{\lambda\leadsto\lambda^{\prime},\bar{v}}$, hence lies in $Z$. As $Z$ is finite and central, passing to a finite extension $\kappa^{\prime}/\kappa(v)$, we see $\sigma_{\lambda\leadsto\lambda^{\prime},v}|_{X_{\kappa^{\prime}}}=\rho_{\lambda\leadsto\lambda^{\prime},v}|_{X_{\kappa^{\prime}}}$. Applying [7, Theorem 1.4.1] again shows that $\sigma_{\lambda,v}|_{X_{\kappa^{\prime}}}=\rho_{\lambda,v}|_{X_{\kappa^{\prime}}}$, so $\sigma_{\lambda,\bar{v}}=\rho_{\lambda,\bar{v}}$, and finally $\sigma=\rho$.
@@
Indeed, by Lemma 3.1 $\rho_{\lambda,\overline{\mathbb{Q}}}$ factors through $\pi_{1}^{t}(X_{\overline{\kappa(v)}},\bar{x})\cong\pi_{1}(X_{\mathcal{O}^{\mathrm{sh}}},\bar{x})$ (this isomorphism follows from tame specialization: see [17, Corollary A.12] and [16, Footnote 3]), and $\rho_{\lambda,x}|_{F^{\mathrm{h}}}$ is unramified, giving the two components of the factorization. The restriction of $\rho_{\lambda}$ to $X_{F^{h}}$ therefore factors through $\pi_{1}(X_{\mathcal{O}^{\mathrm{h}}},\bar{x})$, as this group is isomorphic to $\pi_{1}(X_{\mathcal{O}^{\mathrm{sh}}},\bar{x})\rtimes\pi_{1}(\mathrm{Spec}(\mathcal{O}^{\mathrm{h}}))$, compatibly with the canonical maps from $\pi_{1}(X_{F^{\mathrm{h}}},\bar{x})$: this last point follows by Lemma 3.3, an analogue over a henselian DVR of the exact sequence linking arithmetic and geometric fundamental groups. We are done. [^8] Details: For each $r$, let $Y_{r}\to X[1/N_{\lambda}]$ be the finite étale covering corresponding to the mod $\lambda^{r}$ reduction over a suitable finite extension of $\mathbb{Q}_{\ell}$, and let $\overline{Y}_{r}$ be the normalization of $X^{\prime}=X[1/N_{\lambda}]\cup X_{w}$ in the function field of $Y_{r}$. We must check $\overline{Y}_{r}\to X^{\prime}$ is étale in the fiber above $w$. This is clearly equivalent to checking $\overline{Y}_{r}\times\mathcal{O}_{(w)}\to X^{\prime}_{\mathcal{O}_{(w)}}$ is étale. We are given that $\overline{Y}_{r}\times\mathcal{O}^{\mathrm{h}}\to X^{\prime}_{\mathcal{O}^{\mathrm{h}}}$ is étale. Thus for some finite étale (since $\mathcal{O}_{(w)}$ is a DVR, étale neighborhoods of the special point are explicit) $\mathcal{O}_{(w)}\to\mathcal{O}^{\prime}$, $\overline{Y}_{r}\times\mathcal{O}^{\prime}\to X^{\prime}_{\mathcal{O}^{\prime}}$ is étale (by a limit argument). Thus $\overline{Y}_{r}\to X^{\prime}$ is étale by étale descent.
@@
We imitate [22, Lemma 0BTX]. Surjectivity of $\beta$ is clear: connected objects of $\mathrm{F\acute{E}t}(\mathcal{O})$ are all isomorphic to finite étale subextensions $\mathcal{O}\to\mathcal{O}^{\prime}\subset\mathcal{O}^{\mathrm{sh}}$, and $X_{\mathcal{O}^{\prime}}$ is then connected by assumption. The composite $\beta\circ\alpha$ is trivial; note that for $\mathcal{O}^{\prime}$ as before, $\mathcal{O^{\prime}}\otimes_{\mathcal{O}}\mathcal{O}^{\mathrm{sh}}\equiv\prod_{\mathrm{Gal}(K^{\prime}/K)}\mathcal{O}^{\mathrm{sh}}$ (with $K^{\prime}=\mathrm{Frac}(\mathcal{O}^{\prime})$, $K=\mathrm{Frac}(\mathcal{O})$), just as for fields. To show the image of $\alpha$ is normal, we check that for $U\to X$ finite étale with $U$ connected and such that $U_{\mathcal{O}^{\mathrm{sh}}}\to X_{\mathcal{O}^{\mathrm{sh}}}$ has a section $s$, $U_{\mathcal{O}^{\mathrm{sh}}}$ is a finite coproduct of copies of $X_{\mathcal{O}^{\mathrm{sh}}}$. There is an $\mathcal{O}^{\prime}$ as above such that $s$ is defined over $\mathcal{O}^{\prime}$, and $s(X_{\mathcal{O}^{\prime}})$ is an open connected component of $U_{\mathcal{O}^{\prime}}$ (being a section of an étale separated morphism). Now consider
@@
In particular, this completes the proof in the non-adjoint case of [3, Lemma 5.1], specifically the part deduced from Theorem 5.2 of loc. cit..

### v2

@@ 3. Ramification
In [17, Proposition 2.3], we showed that Zariski-dense $G^{\mathrm{ad}}(\mathbb{C})$-valued representations $\rho$ of $\pi_{1}^{\mathrm{top}}(X(\mathbb{C}))$ in the “superrigid” (loc. cit. Definition 2.2) regime have arithmetic descents with independent of $\ell$ ramification (the same would work for Zariski-dense cohomologically rigid representations with quasi-unipotent local monodromy and finite-order abelianization). Without Zariski-density, and more precisely when the centralizer of $\rho$ is non-trivial, the argument breaks down, and indeed there is no reason that arbitrarily chosen arithmetic descents $\rho_{\lambda}\colon\pi_{1}(X_{F})\to G(\overline{\mathbb{Q}}_{\lambda})$ should have this property, since one can for varying $\lambda$ twist $\rho_{\lambda}$ by homomorphisms $\mathrm{Gal}(\overline{\mathbb{Q}}/F)\to Z_{G}(\rho_{\lambda})(\overline{\mathbb{Q}}_{\lambda})$ with different ramification sets (the basic case to keep in mind, even with $\rho^{\mathrm{ad}}$ Zariski-dense, is when this centralizer is the center of a non-adjoint group $G$). In this section (Lemma 3.2), we show that having one number field specialization with independent of $\ell$ ramification is enough to show that a collection of descents also has this property.
Let $G$ be a connected reductive group; since we will work with sufficiently large coefficients, we assume $G$ is split over $\mathbb{Z}$. In this section we work with a smooth quasi-projective variety $X$ over a number field $F$ whose topological fundamental group $\Gamma=\pi_{1}^{\mathrm{top}}(X(\mathbb{C}),\bar{x})$ (we will specify a base-point later) is $G^{\mathrm{ad}}$-superrigid in the sense of [17, Definition 2.2]: that is, for any algebraically closed field $\Omega$ of characteristic zero and any two Zariski-dense homomorphisms $\rho_{1},\rho_{2}\colon\Gamma\to G^{\mathrm{ad}}(\Omega)$, there is a $\tau\in\operatorname{Aut}(G^{\mathrm{ad}}_{\Omega})$ such that $\tau(\rho_{1})=\rho_{2}$. [^3] Alternatively, throughout this section we could take cohomologically rigid representations with bounded order of quasi-unipotency, without assuming superrigidity properties of $\Gamma$; somewhat longer versions of the arguments then apply using the technique of [9], [16]. We fix once and for all a smooth projective compactification $X\subset\overline{X}$ with $\overline{X}\setminus X$ a strict normal crossings divisor. We further let $N$ be a large enough integer that the setup $(X,\overline{X},\overline{X}\setminus X)$ spreads out to a relative strict normal crossings divisor over $\mathcal{O}_{F}[1/N]$, and we choose some $x\in X(\mathcal{O}_{F^{\prime}}[1/N])$ with $F^{\prime}/F$ a finite extension.
@@
Indeed, this follows from a standard argument (variant of [22, Theorem 4], [9], [16]): the outer action of $\mathrm{Gal}(\overline{\mathbb{Q}}/F)$ on $\pi_{1}(X_{\overline{\mathbb{Q}}},\bar{x})$ induces a discrete action on the finite set $\mathcal{R}_{Z}$, so there is an open subgroup $H\subset\mathrm{Gal}(\overline{\mathbb{Q}}/F)$ such that $H$ fixes each $\rho_{\lambda,\overline{\mathbb{Q}}}$ (up to isomorphism). The adjoint descent $\rho_{\lambda}^{\mathrm{ad}}$ exists uniquely (using Zariski-density: see [17, Lemma 2.1]) over $F(\lambda)^{\mathrm{ad}}:=\overline{\mathbb{Q}}^{H}$. As we vary $\rho$, the $\rho_{\lambda}^{\mathrm{ad}}$ all factor through $G^{\mathrm{ad}}(\mathcal{O})$ for $\mathcal{O}$ the ring of integers in a finite extension of $\mathbb{Q}_{\ell}$, and $G_{Z}(\mathcal{O})\to G^{\mathrm{ad}}(\mathcal{O})$ is on an open subgroup of the source an isomorphism onto its image (an open subgroup of the target). Therefore after restricting to a finite extension $F(\lambda)/F(\lambda)^{\mathrm{ad}}$, each $\rho_{\lambda}^{\mathrm{ad}}$ lifts to
@@
Each $\rho_{\lambda}$ factors through $\pi_{1}(X_{\mathcal{O}_{F(\lambda)}[1/N_{\lambda}]},\bar{x})$ for an integer $N_{\lambda}$ (which we may assume divisible by $\ell$) that depends on $\lambda$ but not on $\rho\in\mathcal{R}_{Z}$. As in [17, Proposition 2.3], set $N^{\prime}=\mathrm{gcd}(N_{\lambda})$, and suppose that $p$ does not divide $N^{\prime}$. We claim that for all $\lambda^{\prime}\nmid p$, $\rho_{\lambda^{\prime},\overline{\mathbb{Q}}}$ factors via the tame specialization map through $\pi_{1}^{t}(X_{\overline{\kappa(v)}},\bar{x}_{v})$ for every $v|p$. Suppose that $p$ divides $N_{\lambda^{\prime}}$ (else the assertion is clear). For some $\lambda\neq\lambda^{\prime}$, $p$ does not divide $N_{\lambda}$. Let $\bar{v}$ be any place of $\overline{\mathbb{Q}}$ above $p$, and let $v=\bar{v}|_{F(\lambda)}$, so for each $\rho\in\mathcal{R}_{Z}$ we have the (tame: see [17, Footnote 3]) restriction to the special fiber
@@
By [8, Theorem 1.4.1], $\rho_{\lambda,v}$ has a (strong) $\lambda^{\prime}$-companion
@@
which clearly lies in $\mathcal{R}_{Z}$. [^5] Had we not assumed superrigidity and instead assumed the elements of $\mathcal{R}_{Z}$ to be cohomologically rigid, this conclusion would still be valid but would require the arguments of [9] and [16]. We may view the elements of $\mathcal{R}_{Z}$ as ¯$\overline{\mathbb{Q}}_{\lambda^{\prime}}$-valued using that they all have representatives factoring through $G(\mathcal{O}_{L})\subset G(\overline{\mathbb{Q}}_{\lambda^{\prime}})$, and we claim that the map $\mathcal{R}_{Z}\to\mathcal{R}_{Z}$, $\rho\mapsto\rho_{\lambda\leadsto\lambda^{\prime},\bar{v}}^{\mathrm{top}}$, is a bijection. Indeed, let $\rho$ and $\sigma$ be elements of $\mathcal{R}_{Z}$ such that $\rho_{\lambda\leadsto\lambda^{\prime},\bar{v}}^{\mathrm{top}}=\sigma_{\lambda\leadsto\lambda^{\prime},\bar{v}}^{\mathrm{top}}$ (equality up to $G$-conjugacy). Then $\rho_{\lambda\leadsto\lambda^{\prime},\bar{v}}=\sigma_{\lambda\leadsto\lambda^{\prime}\bar{v}}$. Let $F_{v}\in\pi_{1}(X_{\kappa(v)},\bar{x}_{v})$ be the $v$-Frobenius induced by the section ($\kappa(v)$-rational point) $x_{v}$. The equality $\rho_{\lambda\leadsto\lambda^{\prime},\bar{v}}(F_{v}\gamma F_{v}^{-1})=\sigma_{\lambda\leadsto\lambda^{\prime},\bar{v}}(F_{v}\gamma F_{v}^{-1})$, for all $\gamma\in\pi_{1}(X_{\overline{\kappa(v)}},\bar{x}_{v})$ implies that $\sigma_{\lambda\leadsto\lambda^{\prime},v}(F_{v})^{-1}\rho_{\lambda\leadsto\lambda^{\prime},v}(F_{v})$ centralizes the image of $\rho_{\lambda\leadsto\lambda^{\prime},\bar{v}}=\sigma_{\lambda\leadsto\lambda^{\prime},\bar{v}}$, hence lies in $Z$. As $Z$ is finite and central, passing to a finite extension $\kappa^{\prime}/\kappa(v)$, we see $\sigma_{\lambda\leadsto\lambda^{\prime},v}|_{X_{\kappa^{\prime}}}=\rho_{\lambda\leadsto\lambda^{\prime},v}|_{X_{\kappa^{\prime}}}$. Applying [8, Theorem 1.4.1] again shows that $\sigma_{\lambda,v}|_{X_{\kappa^{\prime}}}=\rho_{\lambda,v}|_{X_{\kappa^{\prime}}}$, so $\sigma_{\lambda,\bar{v}}=\rho_{\lambda,\bar{v}}$, and finally $\sigma=\rho$.
@@
Indeed, by Lemma 3.1 $\rho_{\lambda,\overline{\mathbb{Q}}}$ factors through $\pi_{1}^{t}(X_{\overline{\kappa(v)}},\bar{x})\cong\pi_{1}(X_{\mathcal{O}^{\mathrm{sh}}},\bar{x})$ (this isomorphism follows from tame specialization: see [18, Corollary A.12] and [17, Footnote 3]), and $\rho_{\lambda,x}|_{F^{\mathrm{h}}}$ is unramified, giving the two components of the factorization. The restriction of $\rho_{\lambda}$ to $X_{F^{h}}$ therefore factors through $\pi_{1}(X_{\mathcal{O}^{\mathrm{h}}},\bar{x})$, as this group is isomorphic to $\pi_{1}(X_{\mathcal{O}^{\mathrm{sh}}},\bar{x})\rtimes\pi_{1}(\mathrm{Spec}(\mathcal{O}^{\mathrm{h}}))$, compatibly with the canonical maps from $\pi_{1}(X_{F^{\mathrm{h}}},\bar{x})$: this last point follows by Lemma 3.3, an analogue over a henselian DVR of the exact sequence linking arithmetic and geometric fundamental groups. We are done. [^7] Details: For each $r$, let $Y_{r}\to X[1/N_{\lambda}]$ be the finite étale covering corresponding to the mod $\lambda^{r}$ reduction over a suitable finite extension of $\mathbb{Q}_{\ell}$, and let $\overline{Y}_{r}$ be the normalization of $X^{\prime}=X[1/N_{\lambda}]\cup X_{w}$ in the function field of $Y_{r}$. We must check $\overline{Y}_{r}\to X^{\prime}$ is étale in the fiber above $w$. This is clearly equivalent to checking $\overline{Y}_{r}\times\mathcal{O}_{(w)}\to X^{\prime}_{\mathcal{O}_{(w)}}$ is étale. We are given that $\overline{Y}_{r}\times\mathcal{O}^{\mathrm{h}}\to X^{\prime}_{\mathcal{O}^{\mathrm{h}}}$ is étale. Thus for some finite étale (since $\mathcal{O}_{(w)}$ is a DVR, étale neighborhoods of the special point are explicit) $\mathcal{O}_{(w)}\to\mathcal{O}^{\prime}$, $\overline{Y}_{r}\times\mathcal{O}^{\prime}\to X^{\prime}_{\mathcal{O}^{\prime}}$ is étale (by a limit argument). Thus $\overline{Y}_{r}\to X^{\prime}$ is étale by étale descent.
@@
We imitate [23, Lemma 0BTX]. Surjectivity of $\beta$ is clear: connected objects of $\mathrm{F\acute{E}t}(\mathcal{O})$ are all isomorphic to finite étale subextensions $\mathcal{O}\to\mathcal{O}^{\prime}\subset\mathcal{O}^{\mathrm{sh}}$, and $X_{\mathcal{O}^{\prime}}$ is then connected by assumption. The composite $\beta\circ\alpha$ is trivial; note that for $\mathcal{O}^{\prime}$ as before, $\mathcal{O^{\prime}}\otimes_{\mathcal{O}}\mathcal{O}^{\mathrm{sh}}\equiv\prod_{\mathrm{Gal}(K^{\prime}/K)}\mathcal{O}^{\mathrm{sh}}$ (with $K^{\prime}=\mathrm{Frac}(\mathcal{O}^{\prime})$, $K=\mathrm{Frac}(\mathcal{O})$), just as for fields. To show the image of $\alpha$ is normal, we check that for $U\to X$ finite étale with $U$ connected and such that $U_{\mathcal{O}^{\mathrm{sh}}}\to X_{\mathcal{O}^{\mathrm{sh}}}$ has a section $s$, $U_{\mathcal{O}^{\mathrm{sh}}}$ is a finite coproduct of copies of $X_{\mathcal{O}^{\mathrm{sh}}}$. There is an $\mathcal{O}^{\prime}$ as above such that $s$ is defined over $\mathcal{O}^{\prime}$, and $s(X_{\mathcal{O}^{\prime}})$ is an open connected component of $U_{\mathcal{O}^{\prime}}$ (being a section of an étale separated morphism). Now consider
@@
In particular, this completes the proof in the non-adjoint case of [4, Lemma 5.1], specifically the part deduced from Theorem 5.2 of loc. cit..